A firm has a corporate bond traded in the secondary market. The maturity of this bond is 5 years and annual coupon interest rate is 12.5%. The bond pays annual coupons and par value is 100$. The market price of this bond is 94.5$. The expected dividend for the next year is 0.75$ per share and the market price of one share is 12$. The corporate tax rate is 20%, the beta of the firm is 1.1, the risk free rate is 9%, and the market risk premium is 6%. The firm targets a debt/equity ratio of 0.5. Estimate the approximate WACC for this firm.
Answer #1
D/A = D/(E+D)
D/A = 0.5/(1+0.5)
=0.3333
Weight of equity = 1-D/A
Weight of equity = 1-0.3333
W(E)=0.6667
Weight of debt = D/A
Weight of debt = 0.3333
W(D)=0.3333
Cost of equity
As per CAPM
Cost of equity = risk-free rate + beta * (Market risk premium)
Cost of equity% = 9 + 1.1 * (6)
Cost of equity% = 15.6
Cost of debt
                  K = N
Bond Price =∑ [(Annual Coupon)/(1 + YTM)^k]     +   Par value/(1 + YTM)^N
                   k=1
                  K =5
94.5 =∑ [(12.5*100/100)/(1 + YTM/100)^k]     +   100/(1 + YTM/100)^5
                   k=1
YTM = 14.106148587
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 14.106148587*(1-0.2)
= 11.2849188696
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=11.28*0.3333+15.6*0.6667
WACC =14.16%